OpenStudy (anonymous):

write an equation in point-slope form for the line that has the given slope and that contains the given point......................slope 4/5 (8,2) choices are y+8=4/5(x+2) y-2=4/5(x-8) y+2=4/5(x+8) y-8=4/5(x-2)

OpenStudy (anonymous):

Point slope form is as follows \[y-a = m(x-b)\] Here you have the slope, m, which is 4/5 And the point (8,2), which is the same as (x,y) You can input these numbers for your answer :)

OpenStudy (anonymous):

@Abs.Zero so would the answer be y-2=4/5(x-8)??

OpenStudy (anonymous):

Yep, that's correct.

OpenStudy (anonymous):

@Abs.Zero thanks

OpenStudy (anonymous):

You're welcome